find (f+g)(x)(f-g)(x)(f x g) and (f divided g)(x) for each f(x)and g(x).
help on this f(x)= 2x
g(x)=-4x+5

Answer by CubeyThePenguin(3113)

(f + g)(x) = f(x) + g(x) = 2x + (-4x + 5) = -2x + 5

(f - g)(x) = f(x) - g(x) = 2x - (-4x + 5) = 6x - 5

(f * g)(x) = f(x) * g(x) = (2x)(-4x + 5) = -8x^2 + 10x

(f/g)(x) = f(x)/g(x) = 2x/(-4x + 5)